# Minibatch Optimal Transport and Perplexity Bound Estimation in Discrete Flow Matching

### v1

@@ 1 Introduction
Modeling distributions of data is a central task in machine learning. In the context of data distributions with continuous support, recent advancements of continuous diffusion and flow models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020; Lipman et al., 2023) have demonstrated impressive capabilities in generating data samples and performing density estimation (Song et al., 2021a, b; Chen et al., 2018) . Rectified flows in particular, have gained attention recently as they enable high quality data generation with very few integration steps. Unfortunately, the application of continuous diffusion and flow models to categorical data distributions remains limited, as they generally underperform compared to autoregressive models in this setting (Chen et al., 2023; Gulrajani & Hashimoto, 2024; Li et al., 2022; Dieleman et al., 2022; Strudel et al., 2022).
To address this, recent research has focused on the development of discrete diffusion (Austin et al., 2021; Campbell et al., 2022; Meng et al., 2022; Lou et al., 2024; Sahoo et al., 2024; Ou et al., 2024) and discrete flow models (Campbell et al., 2024; Gat et al., 2024), which intuitively should be more suitable to model categorical data. In addition, they enable infilling and could potentially speed up generation. In this work, we focus on the discrete flow framework, as this approach increases the design space of possible probability evolutions as compared to discrete diffusion, by enabling the usage of arbitrary couplings and inner dynamics. In addition, discrete and continuous flows share many similarities, facilitating the adoption of results from continuous flow matching to the discrete case. However, despite similarities between the two frameworks, some fundamental differences exist. Such differences are mostly the result of the nonexistence of a discrete flows formulation, in which sample paths are deterministic.
@@ 2.3.1 Probability Flows and Velocities
In discrete flow matching (Gat et al., 2024), the goal is to acquire a flow $p_{t}:[0,1]\times[V]^{L}\rightarrow[0,1]$ constrained by $\sum_{z\in[V]^{L}}p_{t}(z)=1$ that transforms source (reference) distributions $X_{0}\sim p$ to target (data) distributions $X_{1}\sim q$. The flow is completely defined by the choice of a probability velocity $u_{t}:[0,1]\times[V]^{L}\rightarrow\mathbb{R}^{L\times V}$, where $u^{i}_{t}(z)[x^{i}\neq z^{i}]\geq 0$ and $\sum_{x^{i}\in[V]}u^{i}_{t}(z)[x^{i}]=0$, for each position $i$. The update rule of the probability over states when going from time $t$ to $t+\epsilon$ is defined independently for each position in the sequence as follows $p^{i}_{t+\epsilon|t}(x^{i}|x_{t})=\delta_{x^{i}_{t}}(x^{i})+u^{i}_{t}(x^{i},x_{t})$, where we used $u^{i}_{t}(x^{i},z):=u^{i}_{t}(z)[x^{i}]$. Therefore, we can see that as in the framework of Markov chains, the probability over the states in the next step depends solely on the current state, and that $u_{t}$ plays a similar role to a transition-rate matrix $\bar{Q}_{t}$, completely determining the flow. As such, if we learn to approximate the probability velocity $u_{t}(z)$ using a neural network $u_{t}(z;\theta):[0,1]\times[V]^{L}\rightarrow\mathbb{R}^{L\times V}$, we can sample from $p$ and generate data from $q$, using the previous update rule. Before modeling the probability velocity $u_{t}(z)$ however, one must first design an appropriate flow $p_{t}(z)$ that has a suitable, practically learnable corresponding $u_{t}(z)$.
@@ 2.3.2 Conditional Probability Flows
It is clear that this definition of $p_{t}(x|x_{0},x_{1})$ satisfies the conditions $p_{0}(x|x_{0},x_{1})=\delta_{x_{0}}(x)$ and $p_{1}(x|x_{0},x_{1})=\delta_{x_{1}}(x)$. In addition Gat et al. (2024) show that the conditional probability velocity corresponding to Equation (9) is
@@
which in the case of Equations 9/10 implies,
@@
$L=-\mathbb{E}_{t\sim U(0,1)}\mathbb{E}_{x_{0},x_{1}\sim\pi(x_{0},x_{1})}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot|x_{0},x_{1})}\sum_{i=1}^{L}\log p^{i}_{1|t}(x_{0}^{i}|x_{t};\theta).$ (13)
@@ 2.3.3 Source and Target Distributions
As mentioned, points $X_{0}$ and $X_{1}$ that are sampled from a joint distribution $\pi(x,y)$, i.e. $(X_{0},X_{1})\sim\pi(X_{0},X_{1})$, satisfying the marginals constraints $p(x)=\sum_{y\in\mathcal{D}}\pi(x,y)$, $q(y)=\sum_{x\in\mathcal{D}}\pi(x,y)$. As a special case, the training pairs $X_{0}$ and $X_{1}$ can be sampled independently, $(X_{0},X_{1})\sim p(X_{0})q(X_{1})$. Common instantiations of source distribution $p$ are:
@@
$(X_{0},X_{1})=(\mathbf{1}\cdot X_{1}+(1-\mathbf{1})\odot(m,\dots,m),X_{1}),$ (14)
where $X_{1}\sim q(X_{1})$ and $\mathbf{1}\in\{0,1\}^{N}$ is a random variable indicating the conditioning, $\odot$ denotes the entry-wise product, and $\mathbf{1}\in\mathbb{R}^{N}$ is the vector of all ones. They refer to this choice as the conditional coupling (C-coupling).
@@ 3 Minibatch Optimal Transport in the Context of Discrete Flow Matching
One of the main objectives of current research in flow matching is reducing the numbers of steps required for generation. This amounts to simplifying paths/trajectories from the source to the target distribution. An additional benefit to rectifying such paths, is that simplified paths are easier to model by a neural network which empirically has been shown to translate to higher quality models. One can simplify the paths by minimizing the kinetic energy as in the dynamic formulation of optimal transport:
@@
An issue with the categorical dynamic formulation above is that it treats all different tokens equally, while in practice tokens have different degrees of similarity between each other which is reflected in their embeddings. As such, we would like to weight the outflow to each token by some measure of similarity between the donor and recipient of the flow, such that the transition of mass to dissimilar states is particularly penalized. An additional issue is that for large vocabulary sizes, and two sequences sampled from $p(x_{0})$ and $q(x_{1})$ respectively, only a tiny proportion of their positions match exactly. This is in particular the case when the source distribution does not contain information about the target data distribution. This implies that we should not expect substantial improvements by optimizing the Kantorovich expression using minibatches.
@@
where $s(x^{i},x_{t}^{i})\geq 0$ is a similarity metric between two tokens $x^{i}$ and $x_{t}^{i}$, and thus $s(z,z)=0$. Such a similarity metric serves as a penalizing weight for the components of the outflow. In our previous formulation in Equation (17), we had $s(x^{i},x_{t}^{i})=1-\delta_{x_{t}^{i}}(x^{i})$, that is, the discrete metric. Another natural choice is $s(x^{i},x_{t}^{i})=||e_{m}(x^{i})-e_{m}(x_{t}^{i})||^{2}$, which is the $L_{2}$ norm between the embeddings $e_{m}$ of the tokens. Regardless of the similarity metric used, there exists a corresponding Kantorovich formulation, wherein the cost function can be expressed through the similarity metric.
@@
This similarity metric can also be motivated as follows. During training we use the embeddings of the output $o_{i}$ at position $i$ in the last layer of the transformer to get the dot product with respect to the embeddings of each token in the vocabulary, and then try to minimize $-\log{p_{1}}\left(\frac{e^{\langle e_{m}(o_{i}),e_{m}(x_{1}^{i})\rangle}}{\sum_{v}e^{\langle e_{m}(o_{i}),e_{m}(v)\rangle}}\right)=-{\langle e_{m}(o_{i}),e_{m}(x_{1}^{i})\rangle}+A$, where $x_{1}^{i}$ is the token at position $i$ of the pre-perturbed sequence. Therefore, in practice the similarity metric the network uses and adapts the token embeddings for is the dot product $-2{\langle a,b\rangle}=||a-b||^{2}-||a||^{2}-||b||^{2}$.
@@
Flow matching with minibatch OT training involves two optimization procedures: choosing the optimal minibatch coupling, and then training the flow model. These training steps influence one-another as the minibatch coupling changes the flow dynamics, while the token embedding updates during flow matching modify optimal minibatch coupling. While the fact that at each training step the model has only a local view of the flow acts as a stabilizer, we can choose to increase the stability by decoupling the embeddings used in the network and those used in minibatch coupling. One such strategy could for example would be using a moving average of embeddings during minibatch OT coupling.
@@
i) Defining the source distribution $p_{0}=p_{o}$, where $p_{o}(x)=\prod_{i=1}^{L}p^{tok}_{o}(x^{i})$, and $p^{tok}_{o}(x^{i})$ is the empirical probability distribution of tokens in the training set.
@@ 4 An Upper Bound for the Perplexity in Discrete Flow Matching
Finally since the part inside the large brackets is not dependent on $x_{0}$, we can write

### v2

@@ 1 Introduction
Modeling distributions of data is a central task in machine learning. In the context of data distributions with continuous support, recent advancements of continuous diffusion and flow models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020; Lipman et al., 2023) have demonstrated impressive capabilities in generating data samples and performing density estimation (Song et al., 2021a, b; Chen et al., 2018) . Rectified flows, in particular, have gained attention recently as they enable high quality data generation with very few integration steps. Unfortunately, the application of continuous diffusion and flow models to categorical data distributions remains limited, as they generally underperform compared to autoregressive models in this setting (Chen et al., 2023; Gulrajani & Hashimoto, 2024; Li et al., 2022; Dieleman et al., 2022; Strudel et al., 2022).
To address this, recent research has focused on the development of discrete diffusion (Austin et al., 2021; Campbell et al., 2022; Meng et al., 2022; Lou et al., 2024; Sahoo et al., 2024; Ou et al., 2024; Shi et al., 2024) and discrete flow models (Campbell et al., 2024; Gat et al., 2024), which intuitively should be more suitable to model categorical data. In addition, they enable infilling and could potentially speed up generation. In this work, we focus on the discrete flow framework, as this approach increases the design space of possible probability evolutions as compared to discrete diffusion, by enabling the usage of arbitrary couplings and inner dynamics. In addition, discrete and continuous flows share many similarities, facilitating the adoption of results from continuous flow matching to the discrete case. However, despite similarities between the two frameworks, some fundamental differences exist. Such differences are mostly the result of the nonexistence of a discrete flows formulation, in which sample paths are deterministic.
@@ 2.3.1 Probability Flows and Velocities
In discrete flow matching (Gat et al., 2024), the goal is to acquire a flow $p_{t}:[0,1]\times[V]^{L}\rightarrow[0,1]$ constrained by $\sum_{z\in[V]^{L}}p_{t}(z)=1$ that transforms source (reference) distributions $X_{0}\sim p$ to target (data) distributions $X_{1}\sim q$. The flow is completely defined by the choice of a probability velocity $u_{t}:[0,1]\times[V]^{L}\rightarrow\mathbb{R}^{L\times V}$, such that $u_{t}(z)=(u^{1}_{t}(z),\dots,u^{i}_{t}(z),\dots,u^{L}_{t}(z))$ and $u^{i}_{t}:[0,1]\times[V]^{L}\rightarrow\mathbb{R}^{V}$, where $u^{i}_{t}(z)[x^{i}\neq z^{i}]\geq 0$ and $\sum_{x^{i}\in[V]}u^{i}_{t}(z)[x^{i}]=0$, for each $i$. The update rule of the probability over states when going from time $t$ to $t+\epsilon$ is defined independently for each position in the sequence as follows $p^{i}_{t+\epsilon|t}(x^{i}|x_{t})=\delta_{x^{i}_{t}}(x^{i})+u^{i}_{t}(x^{i},x_{t})$, where we used $u^{i}_{t}(x^{i},z):=u^{i}_{t}(z)[x^{i}]$. Therefore, we can see that as in the framework of Markov chains, the probability over the states in the next step depends solely on the current state, and that $u_{t}$ plays a similar role to a transition-rate matrix $\bar{Q}_{t}$, completely determining the flow. As such, if we learn to approximate the probability velocity $u_{t}(z)$ using a neural network $u_{t}(z;\theta):[0,1]\times[V]^{L}\rightarrow\mathbb{R}^{L\times V}$, we can sample from $p$ and generate data from $q$, using the previous update rule. Before modeling the probability velocity $u_{t}(z)$ however, one must first design an appropriate flow $p_{t}(z)$ that has a suitable, practically learnable corresponding $u_{t}(z)$.
@@ 2.3.2 Conditional Probability Flows
It is clear that this definition of $p_{t}(x|x_{0},x_{1})$ satisfies the conditions $p_{0}(x|x_{0},x_{1})=\delta_{x_{0}}(x)$ and $p_{1}(x|x_{0},x_{1})=\delta_{x_{1}}(x)$. In addition, Gat et al. (2024) show that component $i$ of the conditional probability velocity $u_{t}$ corresponding to the flow defined in Equations (8) and (9) is
@@
which in the case of Equations (9) and (10) implies,
@@
$L=-\mathbb{E}_{t\sim U(0,1)}\mathbb{E}_{x_{0},x_{1}\sim\pi(x_{0},x_{1})}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot|x_{0},x_{1})}\sum_{i=1}^{L}\log p^{i}_{1|t}(x_{1}^{i}|x_{t};\theta).$ (13)
@@ 2.3.3 Source and Target Distributions
As mentioned, points $X_{0}$ and $X_{1}$ are sampled from a joint distribution $\pi(x,y)$, i.e. $(X_{0},X_{1})\sim\pi(X_{0},X_{1})$, satisfying the marginals constraints $p(x)=\sum_{y\in\mathcal{D}}\pi(x,y)$, $q(y)=\sum_{x\in\mathcal{D}}\pi(x,y)$. As a special case, the training pairs $X_{0}$ and $X_{1}$ can be sampled independently, $(X_{0},X_{1})\sim p(X_{0})q(X_{1})$. Common instantiations of source distribution $p$ are:
@@
$(X_{0},X_{1})=(\mathbb{I}\cdot X_{1}+(\mathbf{1}-\mathbb{I})\odot(m,\dots,m),X_{1}),$ (14)
where $X_{1}\sim q(X_{1})$ and $\mathbb{I}\in\{0,1\}^{L}$ is a random variable indicating the conditioning, $\odot$ denotes the entry-wise product, and $\mathbf{1}\in\mathbb{R}^{L}$ is the vector of all ones. They refer to this choice as the conditional coupling (C-coupling).
@@ 3 Minibatch Optimal Transport in the Context of Discrete Flow Matching
One of the main objectives of current research in flow matching is to reduce the numbers of steps required for generation. This amounts to simplifying paths/trajectories from the source to the target distribution. An additional benefit to rectifying such paths, is that simplified paths are easier to model by a neural network which empirically has been shown to translate to higher quality models. One can simplify the paths by minimizing the kinetic energy as in the dynamic formulation of optimal transport:
@@
An issue with the categorical dynamic formulation above is that it treats all different tokens equally, while in practice tokens have different degrees of similarity between each other which is reflected in their embeddings. As such, we would like to weight the outflow to each token by some measure of similarity between the donor and recipient of the flow, such that the transition of mass to dissimilar states is particularly penalized. An additional issue is that for large vocabulary sizes, and two sequences sampled from $p(x_{0})$ and $q(x_{1})$ respectively, only a tiny proportion of their positions match exactly. This is especially the case when the source distribution does not contain information about the target data distribution. This implies that we should not expect substantial improvements by optimizing the Kantorovich expression using minibatches.
@@
where $s(x^{i},x_{t}^{i})\geq 0$ is a similarity metric between two tokens $x^{i}$ and $x_{t}^{i}$, and thus $s(z,z)=0$. Such a similarity metric serves as a penalizing weight for the components of the outflow. In our previous formulation in Equation (17), we used $s(x^{i},x_{t}^{i})=1-\delta_{x_{t}^{i}}(x^{i})$, that is, the discrete metric. Another natural choice is $s(x^{i},x_{t}^{i})=||e_{m}(x^{i})-e_{m}(x_{t}^{i})||^{2}$, which is the $L_{2}$ norm between the embeddings $e_{m}$ of the tokens. Regardless of the similarity metric used, there exists a corresponding Kantorovich formulation, wherein the cost function can be expressed through the similarity metric.
@@
Flow matching with minibatch OT training involves two optimization procedures: choosing the optimal minibatch coupling, and then training the flow model. These training steps influence one-another as the minibatch coupling changes the flow dynamics, while the token embedding updates during flow matching modify optimal minibatch coupling when embedding-based similarity metrics are used. While the fact that, at each training step, the model has only a local view of the flow acts as a stabilizer, we can choose to increase the stability by decoupling the embeddings used in the network and those used in minibatch coupling. One such strategy, for example, could be using a moving average of embeddings during minibatch OT coupling.
@@
i) Defining the source distribution $p_{0}=p_{o}$, where $p_{o}(x)=\prod_{i=1}^{L}p^{tok}_{o}(x^{i})$, and $p^{tok}$ is the empirical probability distribution of tokens in the training set.
@@ 4 An Upper Bound for the Perplexity in Discrete Flow Matching
Finally, since the part inside the large brackets is not dependent on $x_{0}$, we can write